OpenStudy (aurorab):

Solve for x: 4/5 x + 4/3 = 2x

OpenStudy (anonymous):

i) multiply both sides by 15 ii) bring both x at one side of eqn. and const on other iii) divide const. by coefficient of x you'll get value of x

OpenStudy (anonymous):

after step 1 eqn will be 12x+20=30x

OpenStudy (anonymous):

2x-4/5x=4/3 (10x-4x)/5=4/3 6x/5=4/3 6x=(4*5)/3 x=20/18 =10/9

OpenStudy (anonymous):

after step 2 eqn is 20= 30x-12x=28 x

OpenStudy (anonymous):

and as madrockz has written you'll get your ans.

OpenStudy (anonymous):

as 20/28=x .. after step 3 which is equal to 10/9=x

OpenStudy (aurorab):

Thank you guys soo much!